OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x,y). 2x + 6y=32 5x - 2y=12

OpenStudy (anonymous):

ok x=32-6y/2 put this in 5x - 2y =12 5(32-6y/2) - 2y = 12 now you can solve it from here ???

OpenStudy (anonymous):

i'll check the answers

OpenStudy (anonymous):

80- 15y -2y=12 80 - 17y =12 y= -68/-17 y =4

OpenStudy (anonymous):

How would that be in (x,y) form?

OpenStudy (anonymous):

wait x= 4 (4,4) is your solution

OpenStudy (anonymous):

in an ordered pair first value is x and second value is y (x,y)